# Quantitative Analysis of Performance Drop in DeepSeek Model Quantization

### v1

@@ 4.3 Main Results
The evaluation results for DeepSeek-R1, DeepSeek-V3 and DeepSeek-R1-distill-Qwen-32B are shown in Table 2, 3, and 4, respectively. We present the official evaluation results reported in (Guo et al., 2025), official deepSeek API [^8] https://api-docs.deepseek.com/ invocation outcomes, and performance metrics of different quantized model variants. For multi-sampling results, we report mean values with corresponding standard deviations (in parentheses). Notably, due to the official DeepSeek V3 API update on March 24, 2025, we substituted it with Tencent’s DeepSeek V3 API  [^9] https://cloud.tencent.com/document/product/1772/115963 in Table 2 to ensure comparability.

### v2

@@ 4.1 Benchmarks
DeepSeek-V3 0324 FP8 (Official API) Q4_K_M (llama.cpp) Q3_K_M (llama.cpp) Q2_K_L (llama.cpp) DQ3_K_M (Ours) Q4_K Q3_K
AIME 2024 57.9 ($\pm$4.34) 53.3 ($\pm$3.10) 54.57 ($\pm$6.14) 31.25 ($\pm$3.04) 57.09 ($\pm$5.16) 59.18 ($\pm$7.91) 52.51 ($\pm$5.29)
MATH 500 93.25 ($\pm$0.91) 93.25 ($\pm$0.47) 92.50 ($\pm$0.96) 85.30 ($\pm$0.68) 93.55 ($\pm$0.25) 93.0 ($\pm$1.06) 91.65 ($\pm$1.34)
GPQA 60.48 ($\pm$1.38) 59.10 ($\pm$1.73) 59.98 ($\pm$0.95) 46.75 ($\pm$0.96) 60.23 ($\pm$1.11) 56.20 ($\pm$2.15) 61.35 ($\pm$2.60)
MBPP 89.03 ($\pm$0.53) 88.63 ($\pm$0.56) 88.10 ($\pm$0.41) 82.93 ($\pm$1.04) 89.50 ($\pm$0.24) 88.43 ($\pm$1.87) 87.78 ($\pm$1.11)
MBPP+ 74.73 ($\pm$0.48) 74.40 ($\pm$0.74) 73.08 ($\pm$0.30) 68.98 ($\pm$1.00) 75.63 ($\pm$0.54) 73.33 ($\pm$2.13) 73.30 ($\pm$1.06)
LiveCodeBench 49.73 ($\pm$1.26) 47.88 ($\pm$1.21) 46.23 ($\pm$0.46) 36.95 ($\pm$0.70) 47.89 ($\pm$0.35) 47.79 ($\pm$1.04) 44.95 ($\pm$0.97)
MMLU 89.08 88.71 88.47 85.59 88.93 88.73 88.57
CMMLU 86.13 86.13 85.28 81.57 85.99 85.96 84.84
C-Eval 89.60 89.10 88 .90 73.60 89.10 89.00 88.50
Average 76.66 75.62 75.24 65.88 76.43 75.74 74.83
Weighted avg. 80.70 80.04 79.56 71.49 80.50 79.81 79.29
Accuracy drop - 1.35% 1.85% 14.66% 0.30% 1.20% 2.39%
Table 4: Quantization results of DeepSeek-V3-0324 on various benchmarks.
@@ 4.3 Main Results
The evaluation results for DeepSeek-R1, DeepSeek-V3 and DeepSeek-R1-distill-Qwen-32B are shown in Table 2, 3, and 5, respectively. We present the official evaluation results reported in (Guo et al., 2025), official deepSeek API [^8] https://api-docs.deepseek.com/ invocation outcomes, and performance metrics of different quantized model variants. For multi-sampling results, we report mean values with corresponding standard deviations (in parentheses). Notably, due to the official DeepSeek V3 API update on March 24, 2025, we substituted it with Tencent’s DeepSeek V3 API  [^9] https://cloud.tencent.com/document/product/1772/115963 in Table 2 to ensure comparability.
@@
DeepSeek-V3-0324
The evaluation results in Table 4 reveal that DeepSeek-V3-0324 maintains strong performance when using our proposed DQ3_K_M method. It achieves near-lossless compression (average drop: 0.30%), outperforming the FP8 baseline on MATH 500 (93.55 vs. 93.25). Extreme 2-bit quantization (Q2_K_L) causes severe degradation (-14.66% average drop), particularly in knowledge-intensive tasks (e.g., C-Eval: 73.60 vs. 89.60). Our method consistently surpasses llama.cpp’s 3-bit variant (Q3_K_M) at the same bit-width and even its 4-bit implementation (Q4_K_M). We also developed fully quantized versions at 3-bit (Q3_K) and 4-bit (Q4_K) precision. DQ3_K_M outperforms both alternatives in average performance metrics. These results demonstrate that DQ3_K_M enables efficient deployment with minimal performance trade-offs.